# Empirical evaluation of normalizing flows in Markov Chain Monte Carlo

### v1

@@ C.2 Hamiltonian Monte Carlo
$\log\alpha_{t}$ $=\log p(x^{\prime}_{t+1})-\log p(x_{t})-0.5\left(r^{(L)\top}_{t}M^{-1}r^{(L)}_{t}-r^{(1)\top}_{t}M^{-1}r^{(1)}_{t}\right).$
@@ C.3 NeuTra MCMC
NeuTra MCMC uses an NF $q(x)$ with the bijection $f$ that maps a data space point $x$ to a latent space point $z=f(x)$. Given a target distribution $p$, it runs an MCMC sampler with the transformed target log density:
$\log\widetilde{p}(x)=\log p(f(x))+\log\left|\det\left(\frac{\partial f(x)}{\partial x}\right)\right|.$
We set the new state as in subsection C.1.
@@ C.4 Jump MCMC and IMH
Every $K$-th step of Jump MCMC with an NF $q(x)$ proposes a new state as in Equation 2 and computes the log acceptance ratio as in Equation 3. We set the new state as in subsection C.1. Other steps are performed with an MCMC sampler. When $K=1$, Jump MCMC reduces to IMH.
@@ D.2 Synthetic non-Gaussian unimodal target distributions
The un-normalized Rosebrock banana log density for an input $x\in\mathbb{R}^{D}$ (with $D$ even and scale $s$) is defined as:
$\log\widetilde{p}(x)=-\sum_{d=1}^{D/2}s(x_{2d-1}^{2}-x_{2d})^{2}+(x_{2d-1}-1)^{2}.$
@@ D.3 Synthetic multimodal target distributions
  • A mixture with three diagonal Gaussian components in 100 dimensions. Component means are $-5,0,$, and $5$, respectively, in all dimensions. Component standard deviations are $0.7$ in all dimensions. Component weights are $1/3$ for all three components.
@@
The un-normalized double well log density for an input $x\in\mathbb{R}^{D}$ is defined as
$\log\widetilde{p}(x)=-\sum_{d=1}^{D}(x^{2}-4)^{2}.$
@@ D.4 Real-world target distributions
We define the real-world target distributions included in our benchmark. We acquire data for likelihood functions from the repository by Magnusson et al. (2024).
@@ D.4.3 Sparse German credit
$\tau$ $=\log(1+\exp(\widetilde{\tau})),\;\lambda=\log(1+\exp(\widetilde{\lambda})),$
@@ D.4.8 Stochastic volatility
Given a parameter vector $(z,\widetilde{\sigma},\widetilde{\mu},\widetilde{\phi}^{\prime})$ with $z\in\mathbb{R}^{3000}$; $\widetilde{\sigma},\widetilde{\mu},\widetilde{\phi}^{\prime}\in\mathbb{R}$ and measurements $y_{i}\in\mathbb{R}$, the 3003D stochastic volatility model is defined as:
@@
$h_{1}$ $=\mu+\sigma z_{1}-\sqrt{1-\phi^{2}},\;h_{i}=\mu+\sigma z_{i}+\phi(h_{i-1}-\mu)\textrm{ for }i>1,$

### v2

@@ C.2 Hamiltonian Monte Carlo
$\log\alpha_{t}$ $=\log p_{X}(x^{\prime}_{t+1})-\log p_{X}(x_{t})-0.5\left(r^{(L)\top}_{t}M^{-1}r^{(L)}_{t}-r^{(1)\top}_{t}M^{-1}r^{(1)}_{t}\right)$.
@@ C.3 NeuTra MCMC
NeuTra MCMC uses an NF $Q$ with the bijection whose inverse map $f^{-1}$ transforms a latent space point $z$ to a target space point $x=f^{-1}(z)$. Given a target distribution $p_{X}(x)$, it runs an MCMC sampler with the transformed target log density:
$\log\widetilde{p}(z)=\log p_{X}(f^{-1}(z))+\log\left|\det\left(\partial_{z}f^{-1}(z)\right)\right|.$
We set the new state as in subsection C.1. This generates a chain of $n$ states. Afterwards, all $n$ latent points $z_{i}$ are transformed to target space samples via $x_{i}=f^{-1}(z_{i})$ for $i=1,\dots,n$.
@@ C.4 Jump MCMC and IMH
Every $K$-th step of Jump MCMC with an NF $Q$ proposes a new state as in Equation 2 and computes the log acceptance ratio as in Equation 3. We set the new state as in subsection C.1. Other steps are performed with an MCMC sampler. When $K=1$, Jump MCMC reduces to IMH.
@@ D.2 Synthetic non-Gaussian unimodal target distributions
The Rosebrock banana log density for an input $x\in\mathbb{R}^{D}$ (with $D$ even and scale $s$) is defined as:
$\log p_{X}(x)=-\sum_{d=1}^{D/2}s(x_{2d-1}^{2}-x_{2d})^{2}+(x_{2d-1}-1)^{2}-C,$
where $C$ is the log of the normalization constant.
@@ D.3 Synthetic multimodal target distributions
  • A mixture with three diagonal Gaussian components in 100 dimensions. Component means are 5,0,-5,0,, and $5$, respectively, in all dimensions. Component standard deviations are $0.7$ in all dimensions. Component weights are $1/3$ for all three components.
@@
The double well log density for an input $x\in\mathbb{R}^{D}$ is defined as
$\log p_{X}(x)=-\sum_{d=1}^{D}(x^{2}-4)^{2}-C,$
where $C$ is the log of the normalization constant.
@@ D.4 Real-world target distributions
We define the real-world target distributions included in our benchmark. We acquire data for likelihood functions from the repository by [39].
@@ D.4.3 Sparse German credit
$\tau$ $),\displaystyle=\log(1+\exp(\widetilde{\tau})),\;\lambda=\log(1+\exp(\widetilde{\lambda}$)),
@@ D.4.8 Stochastic volatility
[…]
@@
$h_{1}$ $=\mu+\sigma z_{1}/\sqrt{1-\phi^{2}},\;h_{i}=\mu+\sigma z_{i}+\phi(h_{i-1}-\mu)\textrm{ for }i>1,$
@@ E.1 Squared bias of the second moment
Ignoring the scaling via the true variance and summarization with the maximum function, $b^{2}$ relates to the classical bias that arises as a component in the bias-variance decomposition of the mean squared error (MSE):
$)\displaystyle\mathrm{MSE}(f(X),f(\hat{X}$)) $=\mathbb{E}[(f(X)-f(\hat{X}))^{2}]$ (17)
$=\underbrace{\mathrm{Var}[f(\hat{X})]}_{\text{Estimator variance}}+\underbrace{(\mathrm{E}[f(X)]-\mathrm{E}[f(\hat{X})])^{2}}_{\text{Squared bias}}+\underbrace{\mathrm{Var}[f(X)],}_{\text{Irreducible error}}$ (18)
where $f$ is a statistical functional, $X$ is the target distribution, and $\hat{X}$ is its approximation, often an empirical distribution based on MCMC samples. The irreducible error term can be ignored within comparisons. Estimator variance decays as $\mathcal{O}(1/n)$ under the Markov Chain central limit theorem [12, Theorem 3], while bias may not vanish as quickly and may thus dominate MSE. For large sample sizes, as in our experiments, we thus expect squared bias to distinguish different methods well.